# Quantifying Misalignment Between Agents: Towards a Sociotechnical Understanding of Alignment

### v2

@@ 6 Case Studies
Shopping Recommender System
Consider an AI-based recommender system employed by a hybrid ”big-box” retailer with both online and brick-and-mortar presence. This system mediates between the retailer and its customers, potentially leading to varying degrees of alignment or misalignment depending on the specific shopping context.
@@
  • Agents: $\mathcal{A}=\{c,R,R^{S}\}$, where $c$ is the customer, $R$ is the retailer, and $R^{S}$ is the recommender system. [^1] Though we only use one instance of each agent here, it is easy to add agents to any category and recalculate misalignment accordingly.
@@
We can now calculate the misalignment for each problem area using our model:
$Pr(0|\mathcal{A},p)=\frac{1}{|\mathcal{A}|^{2}}\sum_{a_{1},a_{2}\in\mathcal{A}}c(g_{a_{1}},g_{a_{2}})\cdot\sqrt{w_{a_{1}}\cdot w_{a_{2}}}$
@@
$Pr(0|\mathcal{A},\mathcal{P})=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}Pr(0|\mathcal{A},p)$
This results in the following scores: $Pr(0|\mathcal{A},p_{f})=0.09$, $Pr(0|\mathcal{A},p_{h})=0.42$, $Pr(0|\mathcal{A},\mathcal{P})=0.26$. This lines up with our intuition that all three agents are much more aligned w/r/t the grocery area than the household area. Overall, this case study demonstrates how our model can capture the nuanced misalignment in e-commerce scenarios, where goals may align in one context (grocery shopping) but conflict in another (impulse shopping), and where recommender system incentives may lead to subtle reward hacking.
Autonomous Vehicle Pre-collision Scenario
We analyze scenario 17 from the CARLA (Car Learning to Act) challenge: ”Obstacle avoidance without prior action.” (Dosovitskiy et al. 2017) This scenario involves an autonomous vehicle encountering an unexpected pedestrian and needing to perform an emergency maneuver.
@@
Goals for both agents in all problem areas are to avoid the described event (e.g., collision, traffic violation). Weights for $a_{v}$ are derived from CARLA penalty coefficients for each event, and estimated for $a_{h}$:
@@
While both agents seemingly hold the same goals (avoiding each event), we assume that pedestrian and AV goals may conflict in how they wish to avoid incident, and simulate this as some level of conflict between $g^{v}_{p}$ and $g^{h}_{p}$ for each $p$.
@@
To obtain population misalignment scores, we simulate a world with 1000 agents that are pedestrians or autonomous vehicles (see Figure 8). As in the goal distribution experiment from Figure 4, a population with opposing goals has the highest misalignment when those goals are most evenly distributed. We note that even when pedestrian and car goals are as conflicted as possible, the high number of problem areas and low weights of most of the problem areas in the autonomous vehicle case study result in low misalignment at almost all times. To confirm that this is indeed an effect of the low weights, we also show what the curves would look like if the weights were all set to 1.
@@ 7 Discussion
Our approach enables modeling misalignment in diverse contexts, from international relations to family disagreements. It captures a variety of real-world phenomena, such as the “strange bedfellows” effect and reward hacking, and allows for misalignment that varies across different populations and problem areas, providing a comprehensive framework for understanding alignment dynamics in diverse contexts. By addressing the fundamental challenge of defining alignment in a world where humans—and not just AI agents—are frequently misaligned, our model affords the analysis of complex scenarios ranging from localized disputes to global conflicts.
Implications for AI Risk. In the context of AI risk analysis, our work contributes to the understanding of existential risks posed by misaligned AI (Avin et al. 2018; Baum et al. 2019; Ord 2020; Bucknall and Dori-Hacohen 2022). By reframing misalignment as primarily a human-centered problem, we align with recent research on AI’s impact on human decision-making and societal structures (Russell 2019; Bak-Coleman et al. 2021; Seger et al. 2020). Our model provides a tool for analyzing potential AI-powered conflicts and their broader implications (Boulanin et al. 2019; Johnson 2019; Lin 2019; Maas, Matteuci, and Cooke 2022), highlighting the complexity of aligning AI with human values.
Limitations. Our model does not address whether an agent’s actions have positive or negative outcomes for the agent itself. The question of how an agent’s goals could be learned remains open, although progress has been made by others in this space (Brown et al. 2021). Due to space constraints, we also do not address how to apply our model to an agent or agents with internally-conflicting goals.
@@ 8 Conclusion
Broader Impacts. Our approach encourages AI safety and alignment researchers to avoid reductionist traps, such as attempting to align AI narrowly with individuals or humanity as a whole, or adopting an overly techno-solutionist mindset. We hope to spark more conversation about the sociotechnical aspects of the alignment problem and the need for interdisciplinary collaboration in addressing these challenges.
Future Work. Future research could extend our model to entities beyond humans and AI, including social constructs like nation-states and corporations (alluded to in the retail case study), as well as biological entities from cellular interactions to ecosystems. This expansion could provide insights into alignment dynamics in diverse complex systems.
@@
Finally, applying our model to existing multi-agent simulations, such as those in MeltingPot (Leibo et al. 2021), DeepMind’s Concordia (Vezhnevets et al. 2023), or other population simulations (Piatti et al. 2024; Foxabbott et al. 2023) could yield valuable insights into the dynamics of misalignment in complex multi-agent systems.
Appendix A AI Risk Analysis
As mentioned in §7, our work contributes to the AI existential risk literature. A 2022 paper presented potential pathways from current and near-term AI to increased x-risk, which does not presuppose AGI (Bucknall and Dori-Hacohen 2022); instead, the authors propose that power struggles such as AI-powered international and state-corporate conflicts may play a large role in increased x-risk (and/or other catastrophic tail risks) due to other, non-AGI risk sources such as nuclear war, runaway climate change, and so forth. Crucially, several papers have argued that current AI is already misaligned with humanity (e.g., Bucknall and Dori-Hacohen 2022; Russell 2019) and also that humanity’s collective sense-making and decision-making capacities are already being compromised and diminished by present-day AI, such as the recommender systems at the core of social media platforms (e.g., Bak-Coleman et al. 2021; Bucknall and Dori-Hacohen 2022; Seger et al. 2020).
By recasting misalignment as first and foremost a human-centered problem, rather than an AI-centric problem, and drawing on existing literature that studies human conflict and contention, our paper ties directly into this line of research that suggests that AI may serve to increase, accelerate and intensify the risks of human conflict—already a thorny and arguably intractable problem even before AI’s advent (Boulanin et al. 2019; Johnson 2019; Lin 2019; Maas, Matteuci, and Cooke 2022). Furthermore, by providing a flexible framework that can be used to account for, analyze and quantify misalignment among a vast array of agents, both human and artificial, our work holds significant promise to advance our field’s understanding of the alignment problem.
Finally, our model encourages AI safety and alignment researchers to avoid the potentially reductionist traps of (a) “narrowly” aligning AI with either individual humans or humanity as a whole by highlighting the challenges in aligning any diverse group of individuals, whether that group includes AI agents or not; and (b) adopting a techno-optimistic and/or techno-positivist mindset that naively supposes that the alignment problem can be solved by technological means alone. We sincerely hope that our paper sparks more conversation in the AI safety and alignment communities about the sociotechnical aspects of the alignment problem, and the need to include a diverse group of researchers and practitioners with expertise in diverse domains far beyond computer science and philosophy departments; and by extension, contributes to improving humanity’s odds of finding realistic and sustainable approaches to reducing x-risk.
Appendix B Entropy
Our maximum-weight, maximum-conflict formulation of misalignment bears an interesting resemblance to entropy in information theory.
For a problem area with $\hat{\mathcal{G}}$ non-zero goals which are mutually exclusive and uniformly distributed, our misalignment measure approaches:
$\lim_{|\mathcal{A}|\to\infty}Pr(0|\mathcal{A},p)=\frac{\hat{\mathcal{G}}-1}{\hat{\mathcal{G}}}$
This can be rewritten in a form that resembles measures of diversity or impurity:
$Pr(0|\mathcal{A},p)\approx 1-\sum_{i=1}^{\hat{\mathcal{G}}}p_{i}^{2}$
where $p_{i}$ is the proportion of agents holding goal $i$.
Interestingly, this formula is equivalent to the Gini impurity, and is the inverse of the Simpson diversity index in ecology. While not directly equivalent to Shannon entropy, it shares some conceptual similarities:
1. Both our misalignment measure and entropy reach their maximum values when the distribution is uniform (i.e., when all goals are equally represented). 2. Both provide a measure of ”disorder” or diversity in a system.
However, there are important distinctions:
1. Shannon entropy is calculated as $-\sum p_{i}\log p_{i}$, which differs from our quadratic form. 2. Our measure has a fixed upper bound of $(\hat{\mathcal{G}}-1)/\hat{\mathcal{G}}$, while Shannon entropy’s upper bound increases with $\hat{\mathcal{G}}$. 3. Our model incorporates additional features like $g_{0}$ (the null goal), pairwise goal conflicts, and importance weights, which are not present in standard entropy calculations.
This connection suggests that groups holding incompatible goals can be viewed as carriers of information, with higher misalignment indicating greater ”disorder” or uncertainty in the system. Entropy or “heat” (in both the metaphoric and literal, thermodynamic senses of the word) would be maximal when misaligned agent groups are each $\frac{|\mathcal{A}_{g}|}{|\mathcal{A}|}$ and identical in size.
Appendix C Assumptions and Restrictions
In the paper as a whole:
@@ Appendix C Assumptions and Restrictions
    The ”no goal” option ($g_{0}$) is not in conflict with any explicit goal.
@@
    Each agent assigns a weight to each problem area, representing its importance.
@@
    Importance ”weights” are non-negative and can sum to more than 1 across problem areas.
@@
    The model assumes that all agents within a population are equally likely to be selected for comparison.
  10. 10.
@@
Extra restrictions in the ”Mutually Exclusive Goals” paragraph:
@@
    Each agent has a single goal for each problem area.
@@
  3. 3.
    The ”no goal” option ($g_{0}$) is not in conflict with any explicit goal.
Appendix D Theorems and Proofs/Derivations
D.1 Misalignment Probability
@@
$Pr(0|\mathcal{A},p):=Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A})\\ \cdot c(g^{1}_{p},g^{2}_{p})$
@@
    The probability of selecting two agents randomly from the population $\mathcal{A}$. Assuming uniform random selection with replacement, this probability is $\frac{1}{|\mathcal{A}|^{2}}$.
@@
    The degree of conflict between the goals of these two agents, given by $c(g^{1}_{p},g^{2}_{p})$.
@@
D.2 Misalignment with Mutually Exclusive Goals
@@
$Pr(0|\mathcal{A},p)=\frac{\sum_{g\in\mathcal{G}}\sum_{g^{\prime}\in\mathcal{G},g^{\prime}<g}2|\mathcal{A}_{g}||\mathcal{A}_{g^{\prime}}|}{|\mathcal{A}|^{2}}$ (1)
@@
    Dividing by $|\mathcal{A}|^{2}$ normalizes by the total number of possible agent pairs.
@@
D.3 Misalignment Upper Bound
@@
Max $\text{Max Misalignment}=\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}$ (2)
@@
    Each goal is held by $\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}$ agents.
@@
    The probability of selecting two agents with different goals is $\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}$.
  3. 3.
    This probability is maximized when goals are mutually exclusive and maximally conflicting.
Therefore, $\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}$ represents the upper bound of misalignment.
D.4 Overall Misalignment Across Problem Areas

### v3

@@ 6 Case Studies
6.1  Shopping Recommender System. Consider an AI-based recommender system employed by a hybrid ”big-box” retailer with both online and brick-and-mortar presence. This system mediates between the retailer and its customers, potentially leading to varying degrees of alignment or misalignment depending on the specific shopping context.
@@
  • Agents: $\mathcal{A}=\{c,R,R^{S}\}$, where $c$ is the customer, $R$ is the retailer, and $R^{S}$ is the recommender system. [^1] Though we only use one instance of each agent here, it is easy to add agents to any category and recalculate accordingly.
@@
We can now calculate misalignment in each problem area:
$Pr(1|\mathcal{A},p)=\frac{\sum_{a_{1},a_{2}\in\mathcal{A}}c(g_{a_{1}},g_{a_{2}})\cdot\sqrt{w_{a_{1}}\cdot w_{a_{2}}}}{|\mathcal{A}|(|\mathcal{A}|-1)}$ (6)
@@
$Pr(1|\mathcal{A},\mathcal{P})=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}Pr(1|\mathcal{A},p)$ (7)
This results in the following scores: $Pr(1|\mathcal{A},p_{f})=0.09$, $Pr(1|\mathcal{A},p_{h})=0.42$, $Pr(1|\mathcal{A},\mathcal{P})=0.26$. This lines up with our intuition that all three agents are much more aligned w/r/t the grocery area than the household area. Overall, this case study demonstrates how our model can capture the nuanced misalignment in e-commerce scenarios, where goals may align in one context (grocery shopping) but conflict in another (impulse shopping), and where recommender system incentives may thus lead to subtle reward hacking.
6.2  Autonomous Vehicle Pre-collision Scenario. We analyze scenario 17 from the CARLA (Car Learning to Act) challenge: ”Obstacle avoidance without prior action.” (Dosovitskiy et al. 2017) This scenario involves an autonomous vehicle encountering an unexpected pedestrian and needing to perform an emergency maneuver.
@@
In Table 3, goals for both agents in all problem areas are to avoid the described event (e.g., collision, traffic violation). Weights for $a_{v}$ are derived from CARLA penalty coefficients for each event, and estimated for $a_{h}$. While both agents prefer to avoid each event, we assume that pedestrian and AV goals may conflict in how they wish to avoid incident, and simulate this as some level of conflict between $g^{v}_{p}$ and $g^{h}_{p}$ for each $p$.
@@
To obtain population misalignment scores, we simulate a world with 1000 agents that are pedestrians or autonomous vehicles (see Figure 8). As in the goal distribution experiment (§5.4), a population with opposing goals has the highest misalignment when those goals are most evenly distributed. We note that even with $c(g^{v}_{p},g^{h}_{p})=1$, the high number of problem areas with low weights results in low misalignment at almost all times. We add curves with maximum weights to confirm that this is indeed an effect of the low weights.
@@ 7 Discussion
Our approach enables modeling misalignment in diverse contexts, from international relations to family disagreements. It captures a variety of real-world phenomena, such as the “strange bedfellows” effect (see §3.4) and reward hacking (see §6.1), and allows for misalignment that varies across different populations and problem areas, providing a comprehensive framework for understanding alignment dynamics in diverse contexts. By addressing the fundamental challenge of defining alignment in a world where humans—and not just AI agents—are frequently misaligned, our model affords the analysis of complex scenarios ranging from localized disputes to global conflicts..
7.1  Implications for AI Risk. In the context of AI risk analysis, our work contributes to the understanding of existential risks posed by misaligned AI (Avin et al. 2018; Baum et al. 2019; Ord 2020; Bucknall and Dori-Hacohen 2022). By reframing misalignment as primarily a human-centered problem, we align with recent research on AI’s impact on human decision-making and societal structures (Russell 2019; Bak-Coleman et al. 2021; Seger et al. 2020). Our model provides a tool for analyzing potential AI-powered conflicts and their broader implications (Boulanin et al. 2019; Johnson 2019; Lin 2019; Maas, Matteuci, and Cooke 2022), highlighting the complexity of aligning AI with human values.
Furthermore, as a framework to measure socio-technical misalignment, our approach could lead to better training procedures that account for value pluralism and more effective regulatory frameworks for ensuring AI systems achieve appropriate levels of alignment across different contexts and populations. See Appendix C for more detail.
7.2  Limitations. Our model does not address whether an agent’s actions have positive or negative outcomes for the agent itself. The question of how an agent’s goals could be learned remains open, although progress has been made by others in this space (Brown et al. 2021). We discuss other assumptions and restrictions in Appendix A.
@@ 8 Conclusion
8.1  Broader Impacts. Our approach encourages AI safety and alignment researchers to avoid reductionist traps, such as attempting to align AI narrowly with individuals or humanity as a whole, or adopting an overly techno-solutionist mindset. We hope to spark conversation about the sociotechnical aspects of the alignment problem and the need for interdisciplinary collaboration in addressing these challenges. We discuss further impacts in Appendix C.
8.2  Future Work. Future research could extend our model to entities beyond humans and AI, including social constructs like nation-states and corporations (alluded to in §6.1), as well as biological entities from cellular interactions to ecosystems. This expansion could provide insights into alignment dynamics in diverse complex systems.
@@
Finally, applying our model to existing multi-agent simulations, such as those in MeltingPot (Leibo et al. 2021), DeepMind’s Concordia (Vezhnevets et al. 2023), and other population simulations (Piatti et al. 2024; Foxabbott et al. 2023) could yield valuable insights into the dynamics of misalignment in complex multi-agent systems.
Appendix A Assumptions and Restrictions
In the paper as a whole, we assume:
@@ Appendix A Assumptions and Restrictions
    Conflict between goals is symmetric:
    $c(g_{x},g_{y})=c(g_{y},g_{x})$
@@
    Each agent assigns a weight to each problem area, representing its importance to that agent.
@@
    Weights are non-negative and can sum to more than 1 across problem areas.
@@
    Each goal considered must be held by at least one agent; otherwise, it should be removed from $\mathcal{G}$. Thus, in each problem area, $|\mathcal{G}|\leq|\mathcal{A}|$.
@@
We also impose the following constraints:
@@
    Each agent holds no more than one goal per problem area.
@@
    A lack of a goal ($g_{0}$) is not in conflict with any explicit goal.
  3. 3.
    All agents are equally likely to be selected.
These constraints mostly hold true in real life, but may have exceptions. For example, human individuals seem to hold contradictory goals in a single problem area. For example, a person may feel conflicted about eating a cookie, because they want to eat less sugar for dietary health reasons, but at the same time they might want to eat the cookie because they would enjoy the experience of eating it.
We can fit this into our model more easily by splitting the problem area of ”eating a cookie” into two – ”consuming sugar” and ”having the experience of eating a cookie”. We hypothesize that for any problem area, if it is possible for one person to hold conflicting preferences as to the resolution of that problem, it is possible to split the problem into two non-conflicting problems, making the situation once again compatible with our model. To address the original problem – whether to eat a cookie, in this case – we propose that a final ”judgment” could be determined by aggregating the sub-problem preferences and scaling according to their respective weights. We leave the testing of our problem-splitting hypothesis and of related aggregation methods for future work.
As for the second constraint (that a lack of a goal is not in conflict with any explicit goal), this might be false in cases where an agent’s goal is related to persuasion, since that implies intending to induce a certain goal or goals in other agents. To apply our model to cases like this, simply allow conflict values between $g_{0}$ and other goals; this is compatible with our model in theory, but excluded from our experiments for simplicity since it is not relevant to most cases. For the case of persuasion in particular, it may be appropriate to add other meta-normative goals (i.e. people’s preferences w.r.t. which goals they hold). These could be goals like “Mind my own business,” “Intervene where there is opportunity to do good,” and “Intervene only if I know the people involved.”
Regarding the third constraint, that all agents are equally likely to be selected, future work could remove this constraint by changing the agent-pair sampling component of the misalignment equation, or by artificially scaling up some parts of the population. However, since the practical value of our model is in aligning AI more fairly, we do not recommend this.
We add an extra restriction in §3.5: Mutually Exclusive Goals:
  1. 1.
@@
Although most real-world cases involve goals with less extreme conflict, the simplified calculation that emerges from this constraint is helpful for deriving the upper bounds of misalignment that emerge in uniformly distributed populations (as in §5.2: Varying Goals, and §5.4: Goal Distribution).
Appendix B Theorems and Proofs/Derivations
B.1 Misalignment Probability
@@
$Pr(1|\mathcal{A},p):=$ $\,\,Pr(a_{1},a_{2}\text{ selected randomly from }\mathcal{A},$ (8)
$a_{1}\neq a_{2})\cdot c(g^{1}_{p},g^{2}_{p})$
@@
    The probability of selecting two different agents randomly from the population $\mathcal{A}$.
@@
    The probability of conflict between the goals of these two agents, given by $c(g^{1}_{p},g^{2}_{p})$.
@@
B.2 Misalignment with Mutually Exclusive Goals
@@
$Pr(1|\mathcal{A},p)=\frac{\sum_{g\in\mathcal{G}}\sum_{g^{\prime}\in\mathcal{G},g^{\prime}<g}2|\mathcal{A}_{g}||\mathcal{A}_{g^{\prime}}|}{|\mathcal{A}|(|\mathcal{A}|-1)}$ (9)
@@
    Dividing by $|\mathcal{A}|(|\mathcal{A}|-1)$ normalizes by the total number of possible agent pairs.
@@
B.3 Misalignment Upper Bound
@@
Max $\text{Max Misalignment}=\frac{|\mathcal{A}|(|\hat{\mathcal{G}}|-1)}{|\hat{\mathcal{G}}|(|\mathcal{A}|-1)}$ (10)
@@
    Each goal $g\in\hat{\mathcal{G}}$ is held by exactly $\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}$ agents
@@
    For any agent $a\in\mathcal{A}$:
    1. (a)
      $a$ shares a goal with $\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}-1$ other agents
    2. (b)
      $a$ is misaligned with $|\mathcal{A}|-\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}$ agents
Therefore, the probability of misalignment between two randomly selected agents is:
$Pr(1|\mathcal{A},p)=\frac{|\mathcal{A}|-\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}}{|\mathcal{A}|-1}$ (11)
Simplifying step by step:
$Pr(1|\mathcal{A},p)$ $=\frac{|\mathcal{A}|-\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}}{|\mathcal{A}|-1}$
$=\frac{\frac{|\mathcal{A}||\hat{\mathcal{G}}|}{|\hat{\mathcal{G}}|}-\frac{|\mathcal{A}|}{|\hat{\mathcal{G}}|}}{|\mathcal{A}|-1}\quad\text{(multiply first term by $\frac{|\hat{\mathcal{G}}|}{|\hat{\mathcal{G}}|}$)}$
[…]
$=\frac{|\mathcal{A}||\hat{\mathcal{G}}|-|\mathcal{A}|}{|\hat{\mathcal{G}}|(|\mathcal{A}|-1)}\quad\text{(multiply both parts by $|\hat{\mathcal{G}}|$)}$
$=\frac{|\mathcal{A}|(|\hat{\mathcal{G}}|-1)}{|\hat{\mathcal{G}}|(|\mathcal{A}|-1)}\quad\text{(factor out $|\mathcal{A}|$ in numerator)}$)
As Figure 4 shows, this probability represents the maximum possible misalignment because any non-uniform distribution would concentrate agents into fewer effective groups. This would increase the probability of randomly selecting aligned agents, thus reducing the overall misalignment probability. This also shows,
Theorem 4: $\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}$ is the exclusive upper bound of misalignment for a population with $|\hat{\mathcal{G}}|$ uniformly distributed non-zero goals.
Proof:
$\lim_{|\mathcal{A}|\to\infty}\frac{|\mathcal{A}|(|\hat{\mathcal{G}}|-1)}{|\hat{\mathcal{G}}|(|\mathcal{A}|-1)}=\frac{|\hat{\mathcal{G}}|-1}{|\hat{\mathcal{G}}|}$ (12)
B.4 Overall Misalignment Across Problem Areas